OpenStudy (anonymous):

John and Tim are looking at the equation the square root of the quantity of 3 times x minus 4 equals square root of x . John says that the solution is extraneous. Tim says the solution is non-extraneous. Is John correct? Is Tim correct? Are they both correct? Justify your response by solving this equation, explaining each step with complete sentences.

OpenStudy (jtvatsim):

The equation is: \[\sqrt{3x-4} = \sqrt{x}\]

OpenStudy (anonymous):

yes

OpenStudy (jtvatsim):

So if you were to try to solve this what would your first step be?

OpenStudy (anonymous):

I have no clue..

OpenStudy (jtvatsim):

Alright, no worries. Is there anything that you don't "like" about the equation. Anything that looks "scary"?

OpenStudy (anonymous):

all of it..

OpenStudy (jtvatsim):

lol, I was afraid you would say that. :) But, don't you agree that it would be much "nicer" if it was 3x - 4 = x instead?

OpenStudy (anonymous):

yeah, cuz then you just add four to both sides and divide by 3

OpenStudy (jtvatsim):

Exactly! That would be great... except that's not the equation we have... But wait! We can grant our own wish if we just square both sides like this: \[(\sqrt{3x-4})^2=(\sqrt{x})^2\] \[3x - 4 = x\]

OpenStudy (anonymous):

so then you have x=(x-4)/3

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

*x+4, sorry

OpenStudy (jtvatsim):

Ok, so we know that x = (x + 4)/3... but that still seems really hard right? Maybe we went the wrong way... But that's okay, let's go back to the last point that was working.

OpenStudy (jtvatsim):

3x - 4 = x

OpenStudy (jtvatsim):

What if I instead subtracted the x from both sides? 2x - 4 = 0 and then added 4 2x = 4 Does that look better? See what you get.

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

so the solution is extraneous then

OpenStudy (jtvatsim):

OK, you are probably right, but how do you know that it is extraneous?

OpenStudy (jtvatsim):

We can always test by plugging it into the equation. If it is true it is "non-extraneous" if it is false it is "extraneous"

OpenStudy (jtvatsim):

It seems to work to me: \[\sqrt{3(2)-4}=\sqrt{2}\] \[\sqrt{6-4} = \sqrt{2}\] \[\sqrt{2}=\sqrt{2}\]

OpenStudy (jtvatsim):

So actually this solution looks "non-extraneous" or "true" to me.